The speeds of vehicles on a highway with speed limit 90 km/h are normally distributed with mean 104 km/h and standard deviation 6 km/h. (Round your answers to two decimal places.)

(a) What is the probability that a randomly chosen vehicle is traveling at a legal speed?
_______ %

(b) If police are instructed to ticket motorists driving 110 km/h or more, what percentage of motorist are targeted?
_______  %
Answer #1

9 = I VOI = 1

90 - 104 a)P(x < 90) = P(254 0%) = P(Z < -2.33) ~ 0.0098

So, 0.98% is the probability that a randomly chosen vehicle is traveling at a legal speed.

110 – 104 b)P(x > 110) = P(Z > 04) = P(Z > 1) = 0.1587

So. 15.87% of motorist are targeted.

Please comment if any doubt. Thank you.